a) integrate (3x+1) √4x^2+12x+5. dx
b) integrate x^2+x+5/(x^2+4)(x+1) dx
1
Expert's answer
2020-05-21T12:32:05-0400

ANSWER a) "\frac { 1 }{ 8 } { \left( 4{ x }^{ 2 }+12x+5 \right) }^{ \frac { 1 }{ 2 } }\left( 8{ x }^{ 2 }+10x-11 \right) +\frac { 7 }{ 2 } \ln { \left| (2x+3)+\sqrt { { (2x+3) }^{ 2 }-4 } \right| } +C"

b)"\ln { \left| x+1 \right| } +\frac { 1 }{ 2 } \arctan { \frac { x }{ 2 } +C }"

EXPLANATION

a)When calculating the integral we use the formulas "\int { \frac { du }{ 2\sqrt { u } } =\sqrt { u } } +C,\quad \int { \frac { dt }{ \sqrt { { t }^{ 2 }-{ a }^{ 2 } } } } =\ln { \left| t+\sqrt { { t }^{ 2 }-{ a }^{ 2 } } \right| } +C" . Denote "u=\sqrt { { t }^{ 2 }{ -a }^{ 2 } } dv=dt" integrating by parts, we calculate the integral "\int { \sqrt { { t }^{ 2 }{ -a }^{ 2 } } dt\quad =t } \sqrt { { t }^{ 2 }{ -a }^{ 2 } } -\int { \frac { { t }^{ 2 } }{ \sqrt { { t }^{ 2 }{ -a }^{ 2 } } } dt } =t\sqrt { { t }^{ 2 }{ -a }^{ 2 } } -\int { \frac { { t }^{ 2 }\ { -a }^{ 2 }+{ \ a }^{ 2 } }{ \sqrt { { t }^{ 2 }{ -a }^{ 2 } } } dt } =" "=t\sqrt { { t }^{ 2 }{ -a }^{ 2 } } -\int { \sqrt { { t }^{ 2 }{ -a }^{ 2 } } dt\ - } { \ a }^{ 2 }\ln { \left| t+\sqrt { { t }^{ 2 }-{ a }^{ 2 } } \right| }" . Add the integral "\int { \sqrt { { t }^{ 2 }{ -a }^{ 2 } } dt\quad \quad }"to both sides of the equality, we obtain "\int { \sqrt { { t }^{ 2 }{ -a }^{ 2 } } dt\quad =\frac { t\sqrt { { t }^{ 2 }{ -a }^{ 2 } } }{ 2 } } -\frac { { \quad a }^{ 2 } }{ 2 } \ln { \left| t+\sqrt { { t }^{ 2 }-{ a }^{ 2 } } \right| } +C" .

"\int { (3x+1)\sqrt { 4{ x }^{ 2 }+12x+5 } dx=\frac { 3 }{ 8 } \int { (8x+12)\sqrt { 4{ x }^{ 2 }+12x+5 } dx } }-" "-\frac { 7 }{ 2 } \int { \ \sqrt { 4{ x }^{ 2 }+12x+5 } dx }=" "=\frac { 3 }{ 8 } \int { \sqrt { 4{ x }^{ 2 }+12x+5 } d(4{ x }^{ 2 }+12x+5) } -\frac { 7 }{ 4 } \int { \sqrt { { (2x+3) }^{ 2 }-4 } d(2x+3) } ="

"=\quad \frac { 1 }{ 4 } { \left( 4{ x }^{ 2 }+12x+5 \right) }^{ \frac { 3 }{ 2 } }-\frac { 7(2x+3)\sqrt { { (2x+3) }^{ 2 }{ -2 }^{ 2 } } }{ 8 } ++\frac { 7 }{ 2 } \ln { \left| (2x+3)+\sqrt { { (2x+3) }^{ 2 }-4 } \right| } +C="

"=\frac { 1 }{ 8 } { \left( 4{ x }^{ 2 }+12x+5 \right) }^{ \frac { 1 }{ 2 } }\left( 8{ x }^{ 2 }+10x-11 \right) +\frac { 7 }{ 2 } \ln { \left| (2x+3)+\sqrt { { (2x+3) }^{ 2 }-4 } \right| }+ +C."

EXPLANATION b)

Using fractional decomposition "\frac { { x }^{ 2 }+x+5 }{ \left( { x }^{ 2 }+4 \right) \left( x+1 \right) \quad } =\frac { { (x }^{ 2 }+4)+(x+1) }{ \left( { x }^{ 2 }+4 \right) \left( x+1 \right) \quad }" "=\frac { { (x }^{ 2 }+4)\quad }{ \left( { x }^{ 2 }+4 \right) \left( x+1 \right) \quad } +\frac { (x+1) }{ \left( { x }^{ 2 }+4 \right) \left( x+1 \right) } =" "=\frac { 1 }{ x+1 } +\frac { 1 }{ { x }^{ 2 }+4. }"

"\int { \frac { { x }^{ 2 }+x+5 }{ \left( { x }^{ 2 }+4 \right) \left( x+1 \right) \quad } dx=\int { \frac { 1 }{ x+1 } dx } } +\int { \frac { 1 }{ { x }^{ 2 }+4. } dx }" "=\ln { \left| x+1 \right| } +\frac { 1 }{ 2 } \arctan { \frac { x }{ 2 } +C }"